# Remote Sensing Image Intelligent Interpretation with the Language-Centered Perspective: Principles, Methods and Challenges

### v1

@@ III-C2 Knowledge Retrieval
In summary, knowledge encoding transforms external knowledge into vector representations that can be understood by models, laying the foundation for subsequent knowledge retrieval. The development of knowledge encoding has gone through three major stages—static encoding, dynamic encoding, and multimodal joint encoding—each enhancing the expressive power of knowledge in different ways. Static encoding methods, such as Word2Vec[93], GloVe[94], and FastText[95], focus on capturing semantic relationships between words. However, they are limited by their inability to dynamically adapt to changes in context. Dynamic encoding methods, such as ELMo[98], BERT[71], and GPT[97], can adjust word embeddings based on contextual information, allowing for more flexible and accurate semantic interpretation. Multimodal joint encoding goes a step further by integrating information from multiple modalities—such as vision and language—thus providing robust support for cross-modal knowledge retrieval and reasoning.
@@
Knowledge retrieval is a crucial step following knowledge encoding, aiming to extract information that is highly relevant to the current task from large-scale knowledge bases or databases. This information is then used to support subsequent reasoning and decision-making processes. With the advancement of technology, knowledge retrieval has evolved from traditional keyword-based search methods to more intelligent approaches, such as static vector-based retrieval and dynamic context-aware reasoning. These modern methods enable models to identify the most valuable and relevant knowledge from massive information repositories, even in complex task scenarios, thereby enhancing task comprehension and decision-making performance.
@@
  • Static Vector Search: Static vector–based knowledge retrieval relies on precomputed knowledge embeddings. As illustrated in Fig.13(a), the system retrieves relevant knowledge by calculating the similarity between the query vector and stored knowledge vectors. Common techniques include TF-IDF[101], k-Nearest Neighbors (k-NN)[102], Approximate Nearest Neighbor (ANN)[103] search, and cosine similarity. TF-IDF[101] evaluates the importance of keywords by measuring their frequency in a document relative to their prevalence across the entire corpus. However, it ignores semantic relationships between words. The k-NN[102] algorithm finds the most similar knowledge vectors by computing distances between the query and the stored vectors. To improve efficiency, ANN[103] uses optimized indexing structures to perform approximate retrieval, reducing computational cost while maintaining acceptable accuracy. Cosine similarity measures the angle between query and document vectors, offering a scale-invariant similarity measure that avoids the influence of vector length. These methods are effective for static vector retrieval but lack adaptability and often struggle with complex semantic understanding tasks, where deeper context and reasoning are required.
    Figure 14: Structural organization of language-centered reasoning and decision-making.
@@
  • Dynamic Contextual Reasoning: Unlike static vector search, dynamic context–aware knowledge retrieval (as illustrated in Fig.13(b)) leverages attention mechanisms[69, 105] to establish relationships between pieces of knowledge within a broader context, significantly improving retrieval precision in complex semantic understanding tasks. The attention mechanism operates by computing attention scores through interactions among queries, keys, and values, enabling the model to identify the most relevant knowledge across the global context. In traditional attention frameworks, the query is matched against all key-value pairs in the knowledge base. However, as the number of knowledge vectors grows, the computational cost of this matching process increases dramatically, leading to long retrieval times. To address this issue, Block-Attention[106] introduces a block-wise strategy, dividing large query sequences into smaller chunks and performing attention calculations within each block independently. This approach significantly reduces computational overhead. For example, with a prompt length of 32K, Block-Attention reduces TTFT-block latency by 98.7%. By optimizing the attention-based retrieval process, this method improves the efficiency of dynamic knowledge retrieval in large-scale knowledge bases.
@@
    In the remote sensing domain, effective text-to-image retrieval is a key challenge for Earth observation tasks. Due to the vast volume of remote sensing imagery and the minimal variance among image embeddings, traditional text retrieval techniques often struggle to match images accurately. To address this, the KTIR[107] (Knowledge-to-Image Retrieval) framework proposes a cross-modal retrieval method. It attributes poor retrieval performance to the lack of differentiation in textual descriptions. KTIR[107] addresses this by converting structured knowledge within remote sensing images into textual expressions, and then applying a Cross-Attention[105] mechanism that uses these expressions as queries for retrieval. This approach significantly improves the retrieval accuracy between remote sensing images and textual descriptions.
@@
In summary, static vector–based retrieval methods perform information retrieval by calculating similarity scores between query and knowledge vectors, but they face significant limitations in capturing contextual variations. To overcome this challenge, dynamic context–aware knowledge retrieval methods leverage attention mechanisms to establish relationships across a broader context, thereby enhancing the accuracy of knowledge retrieval in complex semantic scenarios.
@@
TABLE III: Corresponding representative methods of language-centered reasoning and decision-making.
@@
Task Progress Assessment Behavior Cloning Thought Cloning[110], EBC-LLMAgent[111]
Inverse Reinforcement Learning MA-AIRL[112], Xirl[113]
Reward-Based Liir[114], ExpRD[115]
Policy Learning–Based PPO[116], DQN[117]
Model Predictive Control MPC[118], DMPC[119], L1-NMPC[120]
Decision Making Action-Level Decision Making Hu and Pagilla[121], Namiko[122], PID[123], LQR[124], DeepMPC[125], DeepGait[126], GDQ[127], RobotPilot[128], PIBOT[129]
@@
Tool-Level Decision Making ReACT[90], DFSDT[130], Tool-bench[130], ToolLLaMA[16], Levine et al.[131], LangChain[108], AutoGPT[132], CLIPort[133], Toolformer[134], Re-Invoke[135], WebGPT[136], WebCPM[137]
Figure 15: Illustration of retrospective evaluation methods.
III-D Language-Centered Reasoning and Decision-Making
A core challenge in reasoning and decision-making is how to derive subsequent sensing strategies in a task-driven manner by combining perceptual information, associated knowledge, and available actions. This process involves inferring what to do next based on the current understanding of the task. There are two key problems in this reasoning and decision-making phase: How to assess the gap between external information and task objectives, and How to make decisions based on this gap. The structure of this section is illustrated in Fig.14, and representative methods are summarized in Table III.
III-D1 Task Progress Assessment
Assessing the gap between external information and task objectives is critical for making correct action decisions and satisfying interpretation requirements. In addressing this issue, some studies evaluate the gap by comparing the model’s behavior to expert demonstrations, while others assess it by predicting the potential outcomes of the model’s actions. Based on whether expert demonstrations are used, existing studies can be categorized into two main paradigms: Retrospective Assessment, and Predictive Assessment.
Figure 16: Illustration of predictive evaluation methods.
Retrospective Assessment
Retrospective assessment assumes that the gap between the current state and the task objective can only be determined after task execution is completed. This type of evaluation primarily relies on comparing the model’s behavior with expert demonstrations to assess task discrepancies. There are two main approaches: Behavior Cloning, and Inverse Reinforcement Learning (IRL). Both methods analyze the differences between the model’s behavior and expert behavior, helping to identify deficiencies in the model’s task execution and providing a foundation for further optimization.
  • Behavior Cloning: These methods assess the gap between task objectives and perceptual information by imitating expert demonstration behaviors, as shown in Fig.15(a). The key idea is to treat expert behavior as the learning target, allowing the model to analyze task progress by cloning expert actions. For example, Thought Cloning[110] mimics the human reasoning process, which is not only focusing on the actions themselves, but also modeling the decision-making logic behind them. It quantifies the discrepancy between the model’s behavior and expert goals at both the logical reasoning and execution levels. Similarly, EBC-LLMAgent[111] leverages interpretable mechanisms to link perceptual inputs (e.g., raw data) with expert decision intentions, enabling the model to evaluate how far its behavior deviates from the intended goals. These approaches provide a direct and interpretable pathway for assessing the gap between perceptual information and task objectives by explicitly constructing mappings between behavior and intent.
  • Inverse Reinforcement Learning: Unlike behavior cloning, this class of methods does not directly imitate expert behavior. Instead, it infers the underlying reward function behind expert actions to derive an optimal policy, as shown in Fig.15(b). This approach is particularly suitable for scenarios where expert demonstrations are limited or where designing explicit reward functions is challenging. In vision-language tasks, Inverse Reinforcement Learning (IRL) can be used to infer human decision-making preferences, enabling more reasonable gap assessments and improved model behavior. For example: MA-AIRL[112] employs Generative Adversarial Networks to model reward functions in multi-agent interactive environments, evaluating each agent’s contribution to global goals and quantifying the gap between perceptual information and task objectives. Xirl[113] infers reward functions across different environments to construct a generalizable reward representation for gap assessment. Its core idea is to use visual feature embeddings to achieve cross-environment alignment of perceptual information and use the learned reward function to reflect the deviation between perception and target behavior. Together, these two approaches offer distinct theoretical and practical frameworks for gap evaluation and policy optimization, effectively addressing the challenges posed by complex tasks.
Predictive Assessment
While retrospective assessment methods are effective for evaluating task gaps, they often rely on expert demonstrations, which can be difficult to obtain. To address this limitation, predictive assessment methods have emerged. These approaches aim to evaluate the gap between the current state and task objectives by predicting the model’s own future behavior, thereby determining whether upcoming actions are likely to lead to task success. Within this framework, predictive assessment methods can be further categorized into three types: Reward-based approaches, Policy learning–based approaches, and Model predictive control (MPC)–based approaches. By performing pre-execution predictions, these methods proactively identify potential execution issues and provide guidance to the model, helping it improve performance in real-world tasks.
  • Reward-Based: These methods design reward functions to quantify the gap between external information and task objectives, as shown in Fig.16(a), and use reinforcement learning algorithms to optimize policies that minimize this gap. For example: Liir[114] learns intrinsic rewards for individuals, combining their internal motivations with global objectives. In vision-language tasks, this mechanism enables personalized rewards for sub-goals, allowing for more fine-grained task gap evaluation. ExpRD[115] constructs interpretable reward functions that align reward design with human intuition, helping assess the discrepancy between model behavior and task goals. In such methods, the design of the reward function directly affects both the convergence speed and final performance of the model. Therefore, ensuring the rationality and flexibility of the reward function remains a critical research focus.
  • Policy Learning-Based: These methods optimize decision-making policies directly using policy gradients or value function approximation techniques, as shown in Fig.16(b). The goal is to maximize cumulative rewards, while evaluating task performance by quantifying the gap between the current policy’s return and the optimal policy’s return. Specifically: Value-based algorithms (e.g., DQN[117]) estimate Q-values to assess the potential deviation between current actions and target objectives. Policy-based algorithms (e.g., PPO[116]) optimize the policy distribution directly to narrow the reward gap. A core challenge in this approach is accurately modeling the relationship between the expected return of the current policy and the optimal return. Additionally, balancing exploration and exploitation is essential for improving learning efficiency and policy refinement in complex tasks.
  • Model Predictive Control (MPC): MPC–based methods evaluate the gap between task objectives and perceptual information by dynamically predicting future states, as illustrated in Fig.16(c). These methods emphasize real-time planning during the decision-making process, making them well-suited for scenarios that require continuous control and dynamic adjustment. By constructing a dynamic model, MPC[118] predicts the potential impact of current actions on future states and quantifies the deviation between actions and goals by optimizing a cost function (e.g., reconstruction error or goal matching score), thereby selecting the optimal action. For example: DMPC[119] uses a distributed control framework to decompose a complex global planning task into local optimization problems among multiple agents. It evaluates trade-offs between obstacle avoidance and task objectives, achieving coordinated global planning. L1-NMPC[120] applies adaptive nonlinear control to dynamically optimize the cost function, accurately quantifying the deviation between current actions and target states. It is particularly effective in high-precision control scenarios for gap evaluation and strategy optimization.While these methods demonstrate strong planning capabilities in complex and high-dimensional tasks, they typically involve high computational costs and rely heavily on the accuracy of the dynamic model.
In summary, retrospective assessment relies on expert demonstrations after task completion. It evaluates the model by comparing its behavior with that of experts through behavior cloning or inverse reinforcement learning, helping to identify model deficiencies and offering directions for optimization. However, this approach is limited by the high cost of acquiring expert demonstrations, making it difficult to scale in large-scale, dynamic tasks. In contrast, predictive assessment methods do not require expert demonstrations. Instead, they evaluate the potential impact of model behavior before execution, allowing for early estimation of the likelihood of task success. These methods can be further divided into reward-based, policy learning–based, and model predictive control–based approaches, offering a foundation for real-time strategy adjustments during task execution. Overall, retrospective assessment provides precise behavioral comparisons but depends heavily on expert data, while predictive assessment enhances task adaptability through self-prediction, though its accuracy and stability in prediction remain key challenges to be addressed.
III-D2 Decision Making

### v2

@@ III-C2 Knowledge Retrieval
In summary, knowledge encoding transforms external knowledge into vector representations that can be understood by models, laying the foundation for subsequent knowledge retrieval. The development of knowledge encoding has gone through three major stages—static encoding, dynamic encoding, and multimodal joint encoding—-each enhancing the expressive power of knowledge in different ways. Static encoding methods, such as word2vec[29], GloVe[128], and FastText[67], focus on capturing semantic relationships between words. However, they are limited by their inability to dynamically adapt to changes in context. Dynamic encoding methods, such as ELMo[129], BERT[72], and GPT[135], can adjust word embeddings based on contextual information, allowing more flexible and accurate semantic interpretation. Multimodal joint encoding goes a step further by integrating information from multiple modalities—such as vision and language—thus providing robust support for cross-modal knowledge retrieval and reasoning.
@@
Knowledge retrieval is a crucial step following knowledge encoding. It aims to extract information that is highly relevant to the current task from large-scale knowledge bases or databases. This information is then used to support subsequent reasoning and decision-making processes. With the advancement of technology, knowledge retrieval has evolved from traditional keyword-based search methods to more intelligent approaches, such as static vector-based retrieval and dynamic context-aware reasoning. These modern methods enable models to identify the most valuable and relevant knowledge from massive information repositories, even in complex task scenarios, thereby enhancing task comprehension and decision-making performance.
@@
  • Static Vector Search: Static vector-based knowledge retrieval relies on precomputed knowledge embeddings. As illustrated in Fig.13(a), the system retrieves relevant knowledge by calculating the similarity between the query vector and stored knowledge vectors. Common techniques include TF-IDF[3], k-nearest neighbors (k-NN)[30], approximate nearest neighbor (ANN)[10] search, and cosine similarity. TF-IDF[3] evaluates the importance of keywords by measuring their frequency in a document relative to their prevalence across the entire corpus. However, it ignores semantic relationships between words. The k-NN[30] algorithm finds the most similar knowledge vectors by computing the distances between the query and the stored vectors. To improve efficiency, the ANN[10] uses optimized indexing structures to perform approximate retrieval, thus reducing computational cost while maintaining acceptable accuracy. These methods are effective for static vector retrieval but lack adaptability and often struggle with complex semantic understanding tasks, where deeper context and reasoning are required.
@@
  • Dynamic Contextual Reasoning: Unlike static vector search, dynamic context-aware knowledge retrieval (as illustrated in Fig.13(b)) leverages attention mechanisms[163, 49] to establish relationships between pieces of knowledge within a broader context, significantly improving retrieval precision in complex semantic understanding tasks. The attention mechanism operates by computing attention scores through interactions among queries, keys, and values, enabling the model to identify the most relevant knowledge across the global context. In traditional attention frameworks, the query is matched against all key-value pairs in the knowledge base. However, as the number of knowledge vectors increases, the computational cost of this matching process increases dramatically, leading to long retrieval times. To address this issue, block attention[156] introduces a blockwise strategy, dividing large query sequences into smaller chunks and performing attention calculations within each block independently. This approach significantly reduces the computational overhead. For example, with a prompt length of 32K, block attention reduces the TTFT-block latency by 98.7%. By optimizing the attention-based retrieval process, this method improves the efficiency of dynamic knowledge retrieval in large-scale knowledge bases.
@@
    In the remote sensing domain, effective text-to-image retrieval is a key challenge for Earth observation tasks. Owing to the vast volume of remote sensing imagery and the minimal variance among image embeddings, traditional text retrieval techniques often struggle to match images accurately. To address this shortcoming, the KTIR[115] (knowledge-to-image retrieval) framework proposes a cross-modal retrieval method. It attributes poor retrieval performance to the lack of differentiation in textual descriptions. KTIR[115] addresses this issue by converting structured knowledge within remote sensing images into textual expressions and then applying a cross-attention[49] mechanism that uses these expressions as queries for retrieval. This approach significantly improves the retrieval accuracy between remote sensing images and textual descriptions.
@@
In summary, static vector-based retrieval methods perform information retrieval by calculating similarity scores between query and knowledge vectors, but they face significant limitations in capturing contextual variations. To overcome this challenge, dynamic context-aware knowledge retrieval methods leverage attention mechanisms to establish relationships across a broader context, thereby enhancing the accuracy of knowledge retrieval in complex semantic scenarios.
@@
TABLE IV: Corresponding Representative methods of Language-Centered Reasoning and Decision-Making.
@@
Decision-Making Tool Decision ReACT[190], DFSDT[132], Tool-bench[132], ToolLLaMA[161], Levine et al.[77], LangChain[124], AutoGPT[186], CLIPort[146], Toolformer[139], Re-Invoke[23], WebGPT[119], WebCPM[131]
Action Execution Hu and Pagilla[54], Namiko[137], PID[66], LQR[41], DeepMPC[5], DeepGait[162], GDQ[114], RobotPilot[65], PIBOT[116]
III-D Language-Centered Decision-Making and Execution
Within the closed loop of remote sensing interpretation, the core task of decision-making and execution is to formulate subsequent actionable strategies that advance the interpretation process by synthesizing task requirements, domain knowledge, and the agent’s capability boundaries.
However, the critical bottleneck constraining interpretation quality often stems from perceptual limitations, which manifest at two primary levels:
  • Functional Level: The agent lacks the capacity to directly invoke existing tools or algorithms to satisfy complex interpretation needs;
  • Informational Level: The current observed imagery lacks the essential visual information required to complete the reasoning process.
To address these two levels of challenges, this section explores methods related to tool-based decision-making and action execution. Representative methods are summarized in Table IV.
III-D1 Decision-Making and Execution